OpenStudy (anonymous):

Helpp!! Solve the exponential equation. 98x = 27

OpenStudy (anonymous):

9^8x=27

OpenStudy (bahrom7893):

3^(16x) = 3^3

OpenStudy (bahrom7893):

16x=3

OpenStudy (bahrom7893):

why 24?

OpenStudy (anonymous):

oops go with @bahrom7893

OpenStudy (bahrom7893):

(3^2)^(8x) = 3^3 3^(2*8x) = 3^3 3^(16x) = 3^3 16x=3 x = 3/16

OpenStudy (anonymous):

10^6x=93

OpenStudy (bahrom7893):

u're gonna have to use logs now log10^(6x) = log93 6xlog10 = log93

OpenStudy (bahrom7893):

etc..

OpenStudy (anonymous):

Use a graphing calculator. Solve 5.53x = 805 by graphing. Round to the nearest hundredth.